OpenStudy (anonymous):

A ball is thrown vertically upward from the top of a building 84 feet tall with an initial velocity of 70 feet per second. The distance s (in feet) of the ball from the ground after t seconds is s = 96 + 80t - 16t^2. After how many seconds will the ball touch the ground?

OpenStudy (john_es):

Well I would say the correct formula for "s" with the initial data they give is this, s=84+70t-16t^2 To find the time, you must set s=0. And you will obtain two solutions. The time must be positive so t=5.35 s. Anyways, for the equation you give s = 96 + 80t - 16t^2, the solution is t= 6 s.

OpenStudy (anonymous):

what about the 84 ft and 70 ft/ second?

OpenStudy (john_es):

That's the problem, the correct equation should be s=84+70t-16t^2 (1) That makes use of the data. However they "randomly" (I think) put the other equation and ask a question whose true solution should be the one from (1).